# Studying stellar populations in Omega Centauri with phylogenetics

### v1

@@ B.2 Identification of populations with Gaussian Mixture Models
Figure 12: BIC Scores for the Gaussian Mixture models as a function of number of groups for the Optical sample (upper panel) and the Infrared one (lower panel). The Optical sample prefers 3 Gaussians and the Infrared sample prefers 4 Gaussians.
@@
We define populations by groups that are chemically coherent. To do so, we perform Gaussian Mixture Models (GMM) on the stellar abundance data considering all elements and the element space, because from the PCA exploration we found that all elements contribute to the variance and that abundances are not so correlated with each other. We use the sklearn python package and explore the Bayesian Information Criterion (BIC) scores, which finds a tradeoff between the number of parameters and the model to fit (see Buckley et al. 2024, for extensive discussions about GMM applied to metal-poor stars). BIC scores provide a way to perform a model selection using a penalty proportional to the number of estimated parameters, which help us avoid overfitted solutions. The minimum BIC scores represent the optimal combination of estimated parameters in the GMM model. We thus calculate the BIC scores for different number of groups considering each Gaussian has its own diagonal covariance matrix. The BIC scores are plotted in Fig. 12. The upper panel shows the Optical sample in blue, and the lower panel shows the Infrared sample in red.
@@
The Optical sample supports 3 groups, which corresponds to the lowest BIC scores. The Infrared sample supports 4 groups. The populations found from the metallicity distribution of $\omega$ Cen by Mészáros et al. (2021) (using more stars from the the same Infrared sample as us), Sollima et al. (2005), as well as Alvarez Garay et al. (2024), yield 4 groups. This number is however smaller than the five populations found from the metallicity distribution analyzed by Johnson & Pilachowski (2010), who used a larger version of our Optical sample, or the seven populations found clustering the chemical space using the entire sample of our Infrared sample by Mészáros et al. (2021). This is a further argument against attempting to establish a number of populations in $\omega$ Cen, because it heavily depends on the data, method and the information used. See also further recent discussions on GMM for $\omega$ Cen chemical abundances in Pagnini et al. (2024). In fact the GMM analysis on chemical GALAH data by Buder et al. (2022) shows that the BIC scores and the optimal number of groups have a significant impact when varying the abundance ratios considered for the GMM. In this paper we thus use GMM to see how the various groups might be related to each other and which abundances might be causing the separation of the groups.
@@
Figure 13: Stellar Population Barplot of the Infrared sample. Same as Fig. 5 but using 4 groups which is the optimal number for the Infrared sample from the GMM analysis
@@
Figure 13 shows the Stellar Population Barplot for the Infrared sample, with 4 groups coloured with red, blue, magenta and yellow. As in the case of the Optical sample, the admixture between groups is low, especially for the Group_I1-red. A lightly higher admixture between the Group_I2-blue, Group_I3-magenta and Group_I4-yellow is seen. Following the arguments used for the Optical sample, we can interpret this Stellar Population Barplot by claiming that Group_I1-red might have had a more distinct history than Group_I2-blue, Group_I3-magenta and Group_I4-yellow, which are more related to each other.
@@
Figure 14: Anticorrelations coloured by groups following GMM classification and color scheme of of the GMM models. Each star is colored by the group which has the maximum probability. The right hand panels show the Color Magnitude Diagrams. Top row: optical sample. Bottom row: near infrared sample
@@ B.3 Chemical distributions of the populations
We present a description of the groups general properties. We focus our description on their chemical properties and their stellar parameters only. Regarding the kinematical properties, we investigated the proper motions from Gaia and did not find any particular evidence that some group had different kinematical properties than the other. Regarding the ages, we attempted to estimate ages of the individual stars, but our results could not be validated and thus we could not trust them. Acknowledging that He, C, N and O abundances in $\omega$ Cen are peculiar (Mészáros et al. 2021; Romano et al. 2007; Tailo et al. 2016) and that we do not have He abundances for our stars, we decided not to consider ages in this analysis. Indeed, Marino et al. (2012) and Sollima et al. (2005) extensively discuss on the caveats of using ages to study $\omega$ Cen given these element abundance peculiarities.
The anticorrelations Na-O and Al-O for the Optical sample are plotted in the top panels of Fig. 14. The lower panel shows the anticorrelations Al-Mg and Al-O for the Infrared sample. The right-hand panels are the color magnitude diagrams, this allows a 1-1 comparison between the samples because the parameters come from the same source. This time the stars are coloured by the population with highest frequency for each star, and symbols correspond to if the stars are pure (probability of belonging to the GMM group of more than 90%) or mixed (probability of belonging to the group of less than 90%). Changing this threshold does not change our interpretation, since the admixture of the Stellar Population Barplot is in general very low.
The GMM finds that the populations found are well separated in the anticorrelations. This is expected since already from the PCA we could see that the anticorrelations were dominating the variance of the data (see Fig. 4 and 11). Regarding the Optical sample, G_O2-gray stars have high [Na/Fe] and high [Al/Fe] but low [O/Fe] abundances. G_O3-purple has very high [Na/Fe] but normal [Al/Fe] and [O/Fe]. We note they do not overlap in the red giant branch like the other stars. Pancino et al. (2000) showed that $\omega$ Cen has multiple red giant branches, with the faintest branch being the most metal and calcium rich one. The high [Na/Fe] ratio can be related to a metal-rich progenitor since Na production has a strong dependency of metallicity (Brown et al. 2018). G_O1-orange mix in Na, Al, O and the color-magnitude diagram, belonging to the low-Na, low-Al and high-O side of the anticorrelation. Mixed stars are at the central parts of the diagrams.
@@
Figure 15: Chemical trends coloured by groups following classification and color scheme of Fig. 5 and Fig. 14.
We further explore the chemical abundance trends as a function of metallicity for our populations in Fig. 15. The Optical sample is displayed in the left-hand panels, and the Infrared sample is shown in the right-hand panels. We show the [O/Fe], [Ca/Fe] and the s-process [La/Fe] or [Ce/Fe] abundances.
@@
It is quite interesting to note that there are G_O3-gray stars that are also metal-rich and [Na/Fe] rich, but they follow the red giant branch of the CMD diagram, e.g. are considerably bluer and brighter compared to the purple stars. This can be explained by their difference in oxygen, but this could be also a result of ages, with the gray stars being younger than the purple ones. We are unable to determine individual ages to quantify this difference, given the strong chemical peculiarities of these stars.
The G_I3-magenta stars of the Infrared sample are the most O, Ca and Ce-rich ones, and span a wide range of [Fe/H]. Their location in the color-magnitude diagram is mixed with the rest of the stars, but this could be a selection effect of stellar parameters between these and the Optical sample. G_I4-yellow stars have middle Ca and Ce abundances. G_I1-red, on the other hand, while being those with the highest Al abundances do not have significantly high Ce abundances. G_I2-blue have similar properties to G_O1-orange from the Optical sample, namely low metallicities, O and Ca abundances of the order of 0.4 dex, and low Ce s-process abundances.
@@ Appendix C Infrared sample tree
Figure 17 shows the Infrared tree in the same way as the Optical tree of Fig. 7, which is colored based on the GMM groups illustrated in Fig. 13. The Infrared tree features more sub-branches than the Optical tree, but only one significant ‘caterpillar’ branch, colored with yellow and magenta. However, we categorize these branches according to the populations identified in the Optical tree, which maps the connections of the stars shared between the two datasets (see sections below). The reason for fewer ‘caterpillar’ long branches in this tree compared to the Optical one might be that this dataset uses fewer abundance ratios, making it harder to divide the tree into populations with different histories. We recall that the Stellar Population Barplot of the Infrared sample shows a higher proportion of mixed stars compared to the Optical sample and the GMM model prefers four groups instead of three. It is, of course, also possible that the difference in the number of branches is due to selection effects, because we do not have the same stars in both samples, but our analysis of stars in common does not support this idea.

### v2

@@ B.2 Identification of populations with Gaussian mixture models
Figure 11: BIC Scores for the Gaussian mixture models as a function of number of groups for the Optical sample (upper panel) and the Infrared one (lower panel). The Optical sample prefers three Gaussians and the Infrared sample prefers four Gaussians.
@@
We define populations by groups that are chemically coherent. To do so, we perform Gaussian mixture models (GMM) on the stellar abundance data considering all elements and the element space, because from the PCA exploration we found that all elements contribute to the variance and that abundances are not so correlated with each other. We use the scikit-learn python package (Pedregosa et al. 2011) and explore the Bayesian Information Criterion (BIC) scores, which finds a tradeoff between the number of parameters and the model to fit (see Buckley et al. 2024, for extensive discussions about GMM applied to metal-poor stars). BIC scores provide a way to perform a model selection using a penalty proportional to the number of estimated parameters, which help us avoid overfitted solutions. The minimum BIC scores represent the optimal combination of estimated parameters in the GMM model. We thus calculate the BIC scores for different number of groups considering each Gaussian has its own diagonal covariance matrix. The BIC scores are plotted in Fig. 11. The upper panel shows the Optical sample in blue, and the lower panel shows the Infrared sample in red.
@@
The Optical sample supports three groups, which corresponds to the lowest BIC scores. The Infrared sample supports four groups. The populations found from the metallicity distribution of $\omega$ Cen by Mészáros et al. (2021) (using more stars from the same Infrared sample as us), Sollima et al. (2005), as well as Alvarez Garay et al. (2024), yield four groups. This number is however smaller than the five populations found from the metallicity distribution analyzed by Johnson & Pilachowski (2010), who used a larger version of our Optical sample, or the seven populations found clustering the chemical space using the entire sample of our Infrared sample by Mészáros et al. (2021). This is a further argument against attempting to establish a number of populations in $\omega$ Cen, because it heavily depends on the data, method, and the information used. See also further recent discussions on GMM for $\omega$ Cen chemical abundances in Pagnini et al. (2025). In fact the GMM analysis on chemical GALAH data by Buder et al. (2022) shows that the BIC scores and the optimal number of groups have a significant impact when varying the abundance ratios considered for the GMM. In this paper we thus use GMM to see how the various groups might be related to each other and which abundances might be causing the separation of the groups.
@@
Figure 12: Stellar population bar plot of the Infrared sample. Same as Fig. 4 but using four groups, which is the optimal number for the Infrared sample from the GMM analysis.
@@
Figure 12 shows the stellar population bar plot for the Infrared sample, with four groups colored with red, blue, magenta, and yellow. As in the case of the Optical sample, the admixture between groups is low, especially for the Group_I1-red. A lightly higher admixture between the Group_I2-blue, Group_I3-magenta, and Group_I4-yellow is seen. Following the arguments used for the Optical sample, we can interpret this stellar population bar plot by claiming that Group_I1-red might have had a more distinct history than Group_I2-blue, Group_I3-magenta, and Group_I4-yellow, which are more related to each other.
@@
Figure 13: Anticorrelations colored by groups following GMM classification and color scheme of of the GMM models. Each star is colored by the group which has the maximum probability. The right hand panels show the Color Magnitude Diagrams. Top row: optical sample. Bottom row: near infrared sample. Shaded contours represent the KDE of the datapoints.
@@ B.3 Chemical distributions of the populations
We present a description of the groups general properties. We focus our description on their chemical properties and their stellar parameters only. Regarding the kinematical properties, we investigated the proper motions from Gaia and did not find any particular evidence that some group had different kinematical properties than the other. Regarding the ages, we attempted to estimate ages of the individual stars, but our results could not be validated and thus we could not trust them. Acknowledging that He, C, N, and O abundances in $\omega$ Cen are peculiar (Mészáros et al. 2021; Romano et al. 2007; Tailo et al. 2016) and that we do not have He abundances for our stars, we decided not to consider ages in this analysis. Indeed, Marino et al. (2012) and Sollima et al. (2005) extensively discuss on the caveats of using ages to study $\omega$ Cen given these element abundance peculiarities.
The anticorrelations Na-O and Al-O for the Optical sample are plotted in the top panels of Fig. 13. The lower panel shows the anticorrelations Al-Mg and Al-O for the Infrared sample. The right-hand panels are the color magnitude diagrams, this allows a 1-1 comparison between the samples because the parameters come from the same source. This time the stars are colored by the population with highest frequency for each star, and symbols correspond to if the stars are pure (probability of belonging to the GMM group of more than 90%) or mixed (probability of belonging to the group of less than 90%). Changing this threshold does not change our interpretation, since the admixture of the stellar population bar plot is in general very low.
The GMM finds that the populations found are well separated in the anticorrelations. This is expected since already from the PCA we could see that the anticorrelations were dominating the variance of the data (see Fig. 3 and 10). Regarding the Optical sample, G_O2-gray stars have high [Na/Fe] and high [Al/Fe] but low [O/Fe] abundances. G_O3-purple has very high [Na/Fe] but normal [Al/Fe] and [O/Fe]. We note they do not overlap in the red giant branch like the other stars. Pancino et al. (2000) showed that $\omega$ Cen has multiple red giant branches, with the faintest branch being the most metal and calcium rich one. The high [Na/Fe] ratio can be related to a metal-rich progenitor since Na production has a strong dependency of metallicity (Brown et al. 2018). G_O1-orange mix in Na, Al, O, and the color-magnitude diagram, belonging to the low-Na, low-Al, and high-O side of the anticorrelation. Mixed stars are at the central parts of the diagrams.
@@
Figure 14: Chemical trends colored by groups following classification and color scheme of Fig. 4 and Fig. 13.
We further explore the chemical abundance trends as a function of metallicity for our populations in Fig. 14. The Optical sample is displayed in the left-hand panels, and the Infrared sample is shown in the right-hand panels. We show the [O/Fe], [Ca/Fe], and the s-process [La/Fe] or [Ce/Fe] abundances.
@@
It is quite interesting to note that there are G_O3-gray stars that are also metal-rich and [Na/Fe] rich, but they follow the red giant branch of the CMD diagram, are considerably bluer and brighter compared to the purple stars. This can be explained by their difference in oxygen, but this could be also a result of ages, with the gray stars being younger than the purple ones. We are unable to determine individual ages to quantify this difference, given the strong chemical peculiarities of these stars.
The G_I3-magenta stars of the Infrared sample are the most O, Ca, and Ce-rich ones, and span a wide range of [Fe/H]. Their location in the color-magnitude diagram is mixed with the rest of the stars, but this could be a selection effect of stellar parameters between these and the Optical sample. G_I4-yellow stars have middle Ca and Ce abundances. G_I1-red, on the other hand, while being those with the highest Al abundances do not have significantly high Ce abundances. G_I2-blue have similar properties to G_O1-orange from the Optical sample, namely low metallicities, O and Ca abundances of the order of 0.4 dex, and low Ce s-process abundances.
We plot in the lowest left panel of Fig. 14 the distribution of [Eu/Fe] vs [Fe/H] for the stars colocoloredred by the GMM groups. One can see that the lowest metallicity stars, regardless of the GMM group, have higher [Eu/Fe]. The decreasing trend with [Fe/H] can be noticed, but the scatter is large. This scatter, as well as negative trend with [Fe/H], explains why Eu is important in the PCA (see Fig. 3).
@@ Appendix C Infrared sample tree
Figure 16 shows the Infrared tree in the same way as the Optical tree of Fig. 6, which is colored based on the GMM groups illustrated in Fig. 12. The Infrared tree features more sub-branches than the Optical tree, but only one significant ”caterpillar” branch, colored with yellow and magenta. However, we categorize these branches according to the populations identified in the Optical tree, which maps the connections of the stars shared between the two datasets (see sections below). The reason for fewer ‘caterpillar’ long branches in this tree compared to the Optical one might be that this dataset uses fewer abundance ratios, making it harder to divide the tree into populations with different histories. We recall that the stellar population bar plot of the Infrared sample shows a higher proportion of mixed stars compared to the Optical sample and the GMM model prefers four groups instead of three. It is, of course, also possible that the difference in the number of branches is due to selection effects, because we do not have the same stars in both samples, but our analysis of stars in common does not support this idea.
@@
Appendix D Europium abundances and the connection of $\omega$ Cen with GES
Figure 17: Optical tree. Same as Fig. 6 but showing the trends of [Eu/Fe], [La/Eu], and [Eu/Si] along the branches
In Fig. 17 we plot the same data as in Fig. 6 but the panels show the trends of [Fe/H], [Eu/Fe], [Eu/Si], and [La/Eu] along the branches. In consistency with the scatter plot discussed above, the negative trend of [Eu/Fe] is seen for both, Block-Cen and Child-Cen. The trends of [Eu/Fe] however are not so distinct in both branches one another as in the case of [O/Fe] or [Al/Fe] (see Fig. 6), which can be explained by the fact that [Eu/Fe] becomes important only in the third PC of the PCA, unlike O and Al, which matter in the first and the second PC most. The similarity in the [Eu/Fe] abundances in $\omega$ Cen has already been reported by Johnson & Pilachowski (2010) Block-Cen and could mean that r-process elements might not necessarily have been produced inside $\omega$ Cen in large quantities.
The [La/Eu] and [Eu/Si] trends serve to put in context the possible association of $\omega$ Cen as NSC of the GES. As extensively discussed in the literature, the GES shows significant enhancement of [Eu/Fe] abundances of above 0.5 and even 0.7 dex (Aguado et al. 2021; Matsuno et al. 2021; Carrillo et al. 2022), explained by the r-process produced through a neutron star merger (NSM), which can pollute with Eu an entire dwarf galaxy. The ratio of r/s and r/$\alpha$ elements, such as [Eu/La] and [Eu/Si] can be used to study the importance and the time delay of the r-process pollution, because of the different lifetimes of the different progenitors. In fact, Matsuno et al. (2021) found that GES has a flat trend of [La/Eu] with scatter, in agreement with the trends of [Ba/Eu] of Aguado et al. (2021), consistent with Eu being produced by a NSM event. Because GES has low abundances of $\alpha-$elements compared to the in-situ halo, Monty et al. (2024) uses [Eu/Si] to distinguish between accreted and in-situ stars, showing that accreted stars have significantly higher [Eu/Si] compared to in-situ stars. This includes the globular clusters.
Figure 17 shows that all branches and GMM groups have comparable distributions in [Eu/Fe], and have values that are below the reported ones for GES. The Block-Cen and the Child-Cen branches have increasing [La/Eu] trends with metallicity, suggesting that s-process is important in these populations. [Eu/Si] is low in Block-Cen and has a decreasing trend with metallicity, while for Child-Cen we can not identify a clear trend. The GC-Cen has solar-values for [Eu/Si], and no trend. These abundance trends might rule out the chemical connection of $\omega$ Cen with the GES, but only if we believe that the NSC of the GES has to follow the same chemical trends as the rest of the galaxy, which is not established so far. Indeed, if there was a time-delay to pollute the GES with r-process abundances from a NSM, then it is possible that Block-Cen and GC-Cen formed before this event happened. We recall that $\alpha-$capture element abundances of this population are higher than the typical GES ones. The fact that the Eu abundances do not increase in Child-Cen might also suggest that no other NSM event happened inside $\omega$ Cen.